OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (-4, 6) m = 3/4 y - 6 = 3/4x + 4 y - 6 = 3/4(x - 4) y - 6 = 3/4(x + 4) y + 6 = 3/4(x + 4)

OpenStudy (anonymous):

Are you sure? o:

OpenStudy (anonymous):

oh oops it's y - 6 = 3/4(x + 4)

OpenStudy (the_fizicx99):

y - 6 = 3/4(x + 4)

OpenStudy (anonymous):

Haha, alright, thanks ^-^